Please help:
For f(x) = 5 – 3x and g(x) = x2 – 1, find f(g(x)) and g(f(x)).
thanks

Answer by rmromero(383)


Please help:
For f(x) = 5 – 3x and g(x) = x2 – 1,
find f(g(x))
f(x) = 5 - 3x and g(x) = x^2 - 1
f(g(x)) = 5 - 3(x^2-1)
= 5 - 3x^2 + 3
= 8 - 3x^2

and g(f(x)).
f(x) = 5 - 3x and g(x) = x^2 - 1
g(f(x)) = (5 - 3x)^2 - 1
= 25 - 30x + 9x^2 - 1
= 24 - 30x + 9x^2